OpenStudy (anonymous):

Find the solutions of the equation that are in the interval [0, 2pi). Cos u + cos 2u = 0

OpenStudy (anonymous):

I can do the problem if I can set it up to a single cos function, which is what I have trouble doing.

OpenStudy (anonymous):

Find the solutions of the equation in the interval [0,2∏) cos(X/2)-sin(X)= 0 ------------------ cos(x/2) - 2sin(x/2)cos(x/2) = 0 cos(x/2) = 0 x/2 = 90, 270 x = 180º = pi (540 is 180, and exceeds 2pi) ------------- 1 - 2sin(x/2) = 0 sin(x/2) = 1/2 x/2 = 30º, 150º x = 60º, 300º = pi/3, 5pi/3